### v1

# Data-Free Class-Incremental Gesture Recognition with Prototype-Guided Pseudo Feature Replay

Abstract
Gesture recognition is an important research area in the field of computer vision. Most gesture recognition efforts focus on close-set scenarios, thereby limiting the capacity to effectively handle unseen or novel gestures. We aim to address class-incremental gesture recognition, which entails the ability to accommodate new and previously unseen gestures over time. Specifically, we introduce a Prototype-Guided Pseudo Feature Replay (PGPFR) framework for data-free class-incremental gesture recognition. This framework comprises four components: Pseudo Feature Generation with Batch Prototypes (PFGBP), Variational Prototype Replay (VPR) for old classes, Truncated Cross-Entropy (TCE) for new classes, and Continual Classifier Re-Training (CCRT). To tackle the issue of catastrophic forgetting, the PFGBP dynamically generates a diversity of pseudo features in an online manner, leveraging class prototypes of old classes along with batch class prototypes of new classes. Furthermore, the VPR enforces consistency between the classifier’s weights and the prototypes of old classes, leveraging class prototypes and covariance matrices to enhance robustness and generalization capabilities. The TCE mitigates the impact of domain differences of the classifier caused by pseudo features. Finally, the CCRT training strategy is designed to prevent overfitting to new classes and ensure the stability of features extracted from old classes. Extensive experiments conducted on two widely used gesture recognition datasets, namely SHREC 2017 3D and EgoGesture 3D, demonstrate that our approach outperforms existing state-of-the-art methods by 11.8% and 12.8% in terms of mean global accuracy, respectively. The code is available on https://github.com/sunao-101/PGPFR-3/.
Index Terms:
3D gesture recognition, class-incremental gesture recognition, class-incremental action recognition
I Introduction
Hand gesture recognition is an essential research problem with broad applications in fields such as human-computer interaction, virtual reality, and assistive technology. With the development of deep learning, data-driven gesture recognition methods have made significant progress [1, 2, 3]. However, most of these methods are limited to close-set scenarios and lack the ability to handle unseen or novel gestures. In addition, they often rely on a large amount of annotated data, which is often difficult to obtain in practical applications.
Class-incremental gesture recognition is an important problem within the field of open-set recognition, where the model needs to continually learn and adapt to new gestures. However, catastrophic forgetting poses a significant challenge in this context. One straightforward approach to address this issue is to retain samples from old classes and mix them with samples of new classes during the subsequent incremental training of the model [4, 5]. However, in real-world applications, past data may not always be fully accessible due to privacy and legal constraints associated with user data [6]. Furthermore, the limited memory space of devices poses another challenge, as the number of stored samples increases significantly with the addition of incremental tasks.
Figure 1: Illustration of the proposed data-free class-incremental gesture recognition. The main characteristics of this approach are generating pseudo features in an online fashion during the batch training of new class samples, leveraging prototypes and covariance matrices from old classes.
Data-free class-incremental learning addresses the problems of data privacy concerns and limited memory space. Typical data-free approaches use model inversion to generate pseudo samples [7, 8] or pseudo features [9, 10] of old classes, which would eliminate the need to retain a large generator and only the previously trained network. Recently, Aich et al. [8] introduce a data-free approach for class-incremental hand gesture recognition, leveraging model inversion to generate pseudo samples based on eigenvectors of the covariance matrix and the support vectors of the classifier. However, the samples obtained through model inversion for each task still need to be retained. Notably, this approach stores the inverted samples rather than the original samples, making it highly inefficient. Despite the already significant computational cost of model inversion, the requirement for vast storage to retain these inversion samples escalates as the number of continual tasks increases.
Our objective is to efficiently tackle class-incremental gesture recognition using a data-free approach, as shown in  1. We focus on generating effective pseudo features for old classes, rather than pseudo samples, thereby significantly reducing computational costs. To enhance the representation of past sample distributions, we introduce class prototypes and covariance matrices for old classes. During the batch training of new class samples for each task in continual learning, pseudo features, prototypes, and new features are concurrently fed into the classifier, working together to address the issue of catastrophic forgetting.
To this end, we introduce a framework called Prototype-Guided Pseudo Feature Replay (PGPFR) tailored for class-incremental gesture recognition. The PGPFR comprises four crucial components: Pseudo Feature Generation with Batch Prototypes (PFGBP), Variational Prototype Replay (VPR) for old classes, Truncated Cross-Entropy (TCE) for new classes and Continual Classifier Re-Training (CCRT). To generate a diverse set of pseudo features, the PFGBP combines class prototypes from old classes, features of new classes, and batch-specific class prototypes of new classes, which are computed within each batch of data, utilizing efficient geometric transformations. To strike a balance between overconfidence and uncertainty, temperature sharpening is used to control the steepness of probability predictions of pseudo features. The VPR is built on the assumption that the classifier ought to precisely predict the class associated with the prototype of a particular old class. A plain prototype replay loss is introduced, followed by the VPR that incorporates covariances to enhance robustness. The TCE loss aims to mitigate the influence of domain differences of the classifier [11]. Lastly, we exploit the classifier re-training strategy for class-incremental gesture recognition. It should be noted that the spatial complexity of our model is low, since the pseudo features generated by our approach are discarded immediately after being used in each batch, eliminating the need for their retention.
In summary, the main contributions are listed as follows:
  • We introduce an efficient Prototype-Guided Pseudo Feature Replay (PGPFR) framework for class-incremental gesture recognition. This framework is capable of continually recognizing gestures from new classes while preserving data privacy.
  • We propose a Pseudo Feature Generation with Batch Prototypes (PFGBP) module that utilizes efficient operations and batch prototypes to generate pseudo features and pseudo labels for old classes in an online fashion.
  • We propose Variational Prototype Replay (VPR) which leverages class prototypes and covariance matrices to dynamically update the classifier weights of old classes. Additionally, we introduce Continual Classifier Re-Training (CCRT) strategy to further the problem mitigate catastrophic forgetting.
II Related work
II-A Video-Based Hand Gesture Recognition
Hand gestures primarily consist of various combinations of fingers. Earlier works employed shape matching to measure the dissimilarity among different hand shapes. Deep learning methods learn representations directly from raw data and perform recognition in an end-to-end manner. O Köpüklü et al. [12] introduce a novel hierarchical structure for neural networks, facilitating the efficient deployment of offline-trained CNN architectures using sliding window techniques. The attention mechanism can enable the model to focus on key areas related to gesture recognition in videos or images, thereby improving recognition performance. Consequently, Dhingra et al. [13] propose an end-to-end trainable 3D attention residual neural network Res3ATN, which stacks attention blocks and can adaptively change attention features. In order to better integrate multimodal information. Elboushaki et al. [14] introduce an enhanced motion representation method using iMHI and iDMM, with multimodal fusion strategies at both feature and decision levels. Avola et al. [15] employ an enhanced stacked hourglass network for multitasking, generating 2D heatmaps and hand contours, and predicting 3D hand joint positions and meshes using the MANO layer. Recently, Li et al. [16] propose an analytical framework based on the information bottleneck theory to reduce the interference of gesture irrelevant factors by optimizing feature coding. Although these methods achieve satisfactory performance on different gesture datasets, they rely on the RGB-D video as input. This dependency causes the model susceptible to factors like background noise, illumination variations, and viewpoint changes [8], ultimately leading to a degradation in model performance.
II-B 3D Hand Gesture Recognition
Similar to skeleton-based human action recognition [17, 18, 19], 3D hand gesture recognition refers to the understanding of hand gestures by capturing and analyzing the motion information of human joints. The 3D skeleton sequence is typically utilized as the input due to its inherent advantages in robustness and lightness. Nguyen et al. [20] introduce a novel neural network that learns the SPD matrix from skeletal data and constructs a novel layer using a variant of random gradient descent on the Stiefel manifold. De Smedt et al. [21] argue that hand skeleton data can be used to extract valid hand kinematic descriptors in gesture sequences, statistically and temporally encoded via a Fisher kernel and a multilevel temporal pyramid. Liu et al. [22] argue that gestures can be decoupled into hand pose changes and hand movements and model them separately. Analogously, Guo et al. [23] propose a new edge variation graph to divide each neighborhood of the central node into three groups of physical neighbors, temporal neighbors, and variational neighbors. Graph convolutional networks have proven to be effective in gesture recognition. Song et al. [24] propose the MS-ISTGCN, which uses a spatial and a temporal graph convolutional layer to learn the relationship between hand joints and to capture temporal features, respectively. To prevent spatial information and temporal information from influencing each other to produce fuzzy semantics, Liu et al. [25] propose TD-GCN to capture the spatiotemporal dependence between the skeleton joints by calculating the adjacency matrix of channel dependence and time dependence corresponding to different channels and frames. Transformer also plays an important role in extracting correlations between different nodes and different frames. Zhao et al. [26] propose STST that utilizes a Spatial-Temporal Synchronous Transformer to make the model focus on extracting spatial-temporal correlations simultaneously. Distinct from prior works, we focus on the challenging setting of class-incremental gesture recognition.
II-C Data-Free Class-Incremental Learning
Replay-based incremental learning methods [4, 5] face issues of user data privacy and linear increase in sample storage space with the increasing number of incremental tasks. Knowledge distillation methods [27, 28, 29] use a typical student-teacher architecture to distill the knowledge of past tasks from the teacher model into the student model. To transfer past knowledge to the model under new tasks without retaining past samples or models, data-free approaches [30, 31, 32] utilize the GAN model structure to generate pseudo samples to mitigate catastrophic forgetting. The GAN model needs to be saved throughout the entire CIL process and has much larger computation and storage space compared to the classification model. Recent works use model inversion to generate past samples or features that are mixed with new class samples for model training, such as DeepInversion [7], ABD [11], and BOAT-MI [8]. Specifically, ABD [11] improves DeepInversion [7] by proposing local cross entropy for new class samples and importance-weighted feature distillation for features. BOAT-MI [8] stores the distribution information of past class samples and proposes a combination of SV-based and Proto-based methods to explore pseudo features of old classes in deep feature space. Nevertheless, the computational cost of BOAT-MI is considerable. Another disadvantage of BOAT-MI is the need to retain the pseudo samples obtained by model inversion for each task for subsequent incremental tasks, resulting in a sharp increase in the storage space of inversion samples with the increase of tasks. Our focus is data-free class-incremental gesture recognition without retaining pseudo data, and the objective is to design an efficient pseudo-feature-based approach that fully leverages the advantages of prototypes.
Figure 2: Framework of the proposed data-free class-incremental gesture recognition with Prototype-Guided Pseudo Feature Replay (PGPFR). Backbone is fixed after being trained on $\mathcal{T}_{0}$. In each subsequent class increment task $\mathcal{T}_{i}$, we use class prototypes and covariance matrices to generate pseudo features of old classes through PFGBP, only adjusting classification heads. For the acquisition of new class knowledge, we minimize $\mathcal{L}_{T}$ to learn the decision boundaries of the new class; For preserving knowledge of old classes, we minimize $\mathcal{L}_{P}$ and $\mathcal{L}_{V}$ to strengthen the decision boundaries of old classes.
III Methodology
For Class-Incremental Learning (CIL), the model continuously learns $N$ incremental tasks $\mathcal{T}_{0},\mathcal{T}_{1},...,\mathcal{T}_{i},...,\mathcal{T}_{N}$}. The first task $\mathcal{T}_{0}$ requires $k$ classes, and subsequently, for each incremental task , $d$ new classes $\mathcal{C}_{n}$ are added to the existing classes $\mathcal{C}_{o}$ from all previous old tasks. During each incremental task $\mathcal{T}_{i}$, a specific dataset $\mathcal{D}_{i}$ for $\mathcal{T}_{i}$ is provided for model training. In the context of Data-Free Class-Incremental Learning (DFCIL), the model is invisible to the training set that has already been seen in the past and can only be trained on $\mathcal{D}_{i}$ for $\mathcal{T}_{i}$. The critical issue with DFCIL is catastrophic forgetting, a phenomenon where the decision boundary of the classifier gradually becomes blurred as more incremental tasks are introduced, ultimately resulting in a substantial decline in model performance.
We propose an effective Prototype-Guided Pseudo Feature Replay (PGPFR) framework. On the one hand, we strengthen the decision boundaries of old classes by introducing Pseudo Feature Generation with Batch Prototypes (PFGBP) and Variational Prototype Replay (VPR) for old classes. On the other hand, we employ Truncated Cross-Entropy (TCE) to learn decision boundaries for new categories, alongside the Continual Classifier Re-Training (CCRT) strategy, which addresses the challenges of overfitting to instances of new classes and underfitting to those of old classes.
For convenience, we divide the model into two parts: a backbone network $f(\cdot)$ and a classifier $\varphi(\cdot)$. DG-STA [33] is employed as the backbone to effectively extract 3D gesture features, and a fully connected layer serves as the linear classifier for gesture recognition. The overall architecture of the proposed method is illustrated in Figure 2, and details are as follows.
III-A Pseudo Feature Generation with Batch Prototypes
To strengthen the decision boundary of the old classes, Pseudo Feature Generation with Batch Prototypes (PFGBP) is designed to generate pseudo features of old classes. Inspired by [9], we obtain pseudo features by performing simple geometric translations of prototypes of old classes in the feature space. The pseudo feature generated by the PFGBP is represented as:
$\hat{f}(x_{p})=f(x_{n})+\mu_{p}-\hat{\mu}_{n},$ (1)
where $n$ and $p$ are indexes of new class and previous classes, respectively, $x_{n}$ is a sample of the new class, $\mu_{p}$ is the prototype of the $p$-th class, $\hat{\mu}_{n}$ is the estimated prototype for the $n$-th class in a batch of data, $\hat{f}(x_{p})$ is the generated pseudo feature corresponding to the $p$-th class.
The inclusion of all new class samples in the input guarantees that pseudo features and new class samples exhibit a comparable distribution, thereby addressing the issue of data imbalance. Simple geometric translations cause the pseudo features to reside in the feature space close to the prototype of the past class, indicating the validity of pseudo features.
The pseudo labels of old classes corresponding to the generated pseudo features are selected as follows. For each batch of new class samples, we first calculate the batch class prototype, denoted as $\hat{\mu}_{n}$. Then, the similarities between $\hat{\mu}_{n}$ and prototypes of old classes are computed. The most similar past class prototype is chosen as $\mu_{p}$, and the corresponding class serves as the pseudo label $\hat{y}$. The formula for assigning labels to pseudo features can be expressed as:
$\hat{y}=\mathrm{argmax}_{0\leq p\leq\mathcal{N}_{o}}\mathrm{sim}(\hat{\mu}_{n},\mu_{p}),$ (2)
where $\mathcal{N}_{o}$ is a number of old classes, and $\mathrm{sim}(\cdot)$ denotes the cosine similarity function.
It should be noted that we utilize batch class prototypes instead of the class prototypes derived from all samples within a given class. By leveraging batch class prototypes, we can generate a more diverse set of pseudo features. Consider the scenario that $\mathcal{T}_{n}$ comprises only a single new class; in such a case, the pseudo label $\hat{y}$ and its corresponding $\mu_{p}$ would remain fixed to the same old category throughout the $n$-th task. This constraint limits the diversity of pseudo features and ultimately prevents the model from effectively revisiting knowledge from other older classes. Compared with the original prototype, the batch class prototype could better represent the characteristics of the individual data in the batch. Namely, using all samples to compute the prototype may capture some noise or unnecessary details within the class, while computing the prototype using only a batch can reduce this risk, thereby generating a more representative prototype. The specific generation process of pseudo features and pseudo labels is outlined in Algorithm 1.
The pseudo features $\hat{f}(x_{p})$ are concatenated with the features of the new class samples $f(x_{n})$ to form $\tilde{Z}=[\hat{f}(x_{p}),f(x_{n})]$. The corresponding labels $\hat{y}$ and y are also concatenated to form $\tilde{Y}=[\hat{y},y]$. Finally, $\tilde{Z}$ and $\tilde{Y}$ are fed into the classifier, and the cross entropy between the predicted distributions and the labels are:
$\mathcal{L}_{P}=\sum_{z_{k}\in\tilde{Z},y_{k}\in\tilde{Y}}l(z_{k},y_{k}),$ (3)
where $l(z,y)$ is defined as
$l(z,y)=\begin{cases}-\sum_{i=1}^{\mathcal{N}_{o}}y_{i}\log(\mathrm{softmax}(\varphi(z)/R))\text{, if }z\in\hat{f}(x_{p})\\ -\sum_{i=1}^{\mathcal{N}_{o}}y_{i}\log(\mathrm{softmax}(\varphi(z)))\text{, else }z\in f(x_{n}),\end{cases}$ (4)
where $R$ is a temperature parameter to sharpen the predictive distribution of pseudo features, y is a one-hot vector corresponding to a real label or pseudo label and $\mathrm{softmax}(\cdot)$ is an activation function that normalizes classified logits.
It should be noted that temperature sharpening adjusts the probability distribution output by the softmax function through the parameter $R$. When $R<1$, the distribution becomes sharper, meaning the probabilities for the most likely classes increase while the probabilities for less likely classes decrease. This strategy could reduce the model’s sensitivity to uncertain classes, mitigating the negative impact of noise on training.
Algorithm 1 The generation process of pseudo features and pseudo labels
# O: the number of past class prototypes
# B: represents the number of batches
# T: Transposition operation
Input: a batch of samples $x_{n}={x}_{1:B}$, the feature extraction network $f(\cdot)$ in the current task, prototype set $\mathcal{P}_{\text{old}}={\{\mu_{i}\}_{1}^{O}}$.
Output: a set of pseudo features $\mathcal{F}_{\text{pseudo}}$ and a set of pseudo labels $\mathcal{L}_{\text{pseudo}}$.
1:  # Initialize the set of pseudo features and pseudo labels
2:  $\mathcal{F}_{\text{pseudo}}\leftarrow\{\},\mathcal{L}_{\text{pseudo}}\leftarrow\{\}$
3:  # Utilize features the extraction network $f(\cdot)$ extract features $f(x_{n})$ from $x_{n}$
[…]
5:  # Calculate the means of $f(x_{n})$ corresponding to labels to obtain Batch Class Prototypes $\hat{\mu}_{n}$
6:  $f(x_{n})\overset{Mean}{\to}\hat{\mu}_{n}$
7:  # Calculate the similarity matrix $C$ between $\hat{\mu}_{n}$ and $\mathcal{P}_{\text{old}}$
8:  $C\leftarrow\hat{\mu}_{n}\cdot\mathcal{P}_{\text{old}}^{T}$
9:  # Select the old category p that is most similar to $\hat{\mu}_{n}$
10:  $p\leftarrow\arg\max_{j}\,C_{ij}$
11:  # Generate pseudo features using Eq.(1)
12:  $\hat{f}(x_{p})\leftarrow f(x_{n})+\mu_{p}-\hat{\mu}_{n}$
13:  # Merge $\hat{f}(x_{p})$ and $p$ into $\mathcal{F}_{\text{pseudo}}$ and $\mathcal{L}_{\text{pseudo}}$
14:  $\mathcal{F}_{\text{pseudo}}\leftarrow\mathcal{F}_{\text{pseudo}}\cup\{\hat{f}(x_{p})\}$
15:  $\mathcal{L}_{\text{pseudo}}\leftarrow\mathcal{L}_{\text{pseudo}}\cup\{p\}$
III-B Variational Prototype Replay for Old Classes
For old classes, the prototype of a particular class should be as close as possible to the classifier’s weight for that class, while being as far away as possible from the classifier’s weights for other classes. In other words, the classifier should accurately predict the class for the prototype of a given old class. We define prototype replay loss to enable the classifier to better distinguish the differences between different prototypes of old classes. The loss is formulated as:
$\mathcal{L}_{Proto}=\frac{1}{\mathcal{N}_{o}}\sum^{\mathcal{N}_{o}}_{k=1}-\log(\frac{e^{w^{T}_{k}\mu_{k}+b_{k}}}{\sum^{\mathcal{N}_{o}}_{c=1}e^{w^{T}_{c}\mu_{k}+b_{c}}}),$ (5)
where $\mathcal{N}_{o}$ is a number of old classes, $w_{k}$ is the classification weight for the $k$-th class in the fully connected layer, $b_{k}$ is the corresponding bias vector, and $\mu_{k}$ is the prototype of the $k$-th class.
The original prototype lacks a measure of sample distributions of the given class. Samples closer to class prototypes in the feature space are more likely to be classified correctly, while those farther away may be misclassified, which leads to a single solution of the intra-class features. We employ the class prototypes and covariance matrices, which characterizes each class as a distribution rather than a singular point in the latent space. To enhance the robustness and generalization, we incorporate covariances and penalize the predictions of class prototypes with a regularization term in the denominator. The Variational Prototype Replay (VPR) loss is formulated as follows:
$\mathcal{L}_{V}=\frac{1}{\mathcal{N}_{o}}\sum^{\mathcal{N}_{o}}_{k=1}-\log(\frac{e^{w^{T}_{k}\mu_{k}+b_{k}}}{\sum^{\mathcal{N}_{o}}_{c=1}e^{w^{T}_{c}\mu_{k}+b_{c}+\gamma(w_{c}-w_{k})^{T}C_{k}(w_{c}-w_{k})}}),$ (6)
where $\mathcal{N}_{o}$ is the number of old classes, $\mu_{k}$ is the class prototype of the $k$-th class, $C_{k}$ is the covariance matrix of samples of the $k$-th class, and $\gamma$ is a hyperparameter.
As the covariance matrix is symmetric, the penalty term in the denominator is positive. To minimize the loss $\mathcal{L}_{V}$, the optimizer would increase the numerator which represents the likelihood of the classifier accurately classifying the class prototypes, thereby improving the consistency between the classifier and the prototypes.
III-C Truncated Cross-Entropy for New Classes
As mentioned in [11], there exists a domain difference between real data and pseudo features generated by model inversion. When pseudo features and real data are mixed for model training, the model tends to classify the mixed samples as either real or synthetic, rather than discriminating them based on the semantic information encoded within the features. To mitigate the influence of domain differences of the classifier, we introduce an additional classifier solely dedicated to the new classes. This classifier, termed Truncated Cross-entropy for New Classes (TCE), is formulated as follows:
$\mathcal{L}_{T}=\sum_{x\in\mathcal{D}^{n}}\sum_{y\in\mathcal{T}_{n}}\ -y\mathrm{log}(\mathrm{softmax}(\varphi^{n}(f(x))))),$ (7)
where $\varphi^{n}(\cdot)$ is the classification header of the $n$-th task, $f(\cdot)$ is the feature extractor.
III-D Continual Classifier Re-Training
The overall training objective loss is:
$\mathcal{L}=\mathcal{L}_{P}+\mathcal{L}_{V}+\mathcal{L}_{T}.$ (8)
During training, we design Continual Classifier Re-Training (CCRT) strategy, which is inspired by classifier re-training [34, 35] in long-tail learning. However, in class incremental learning, this training strategy is rarely studied. Specifically, we fix the backbone after training it on the first task and keep the backbone parameters unchanged throughout subsequent incremental tasks. For each subsequent incremental task, we only re-train the classifier to adjust the decision boundary. This strategy guarantees the stability of extracted features, preserving a consistent feature space for pseudo feature generation. Fixing the backbone allows for the acquisition of a more generalized feature representation, which aids the classifier in adapting to new tasks. It also avoids overfitting to new classes and retains feature information from old classes, helping to mitigate catastrophic forgetting. The training procedure is illustrated in the supplementary material.
IV Experiment
IV-A Dataset
The SHREC 2017 3D dataset, derived from Online-DHG, includes 14 gesture sequences performed by 28 participants, totaling 2800 sequences. It captures both coarse and fine gestures using 22 hand joints in 2D and 3D space, Each of which is represented by a single coordinate. The skeletal structure is represented by the coordinates of 22 joints in both 2D depth image space and 3D world space, constituting a complete hand representation. These 22 joint points include 1 palm, 1 wrist, and 4 joints per finger, totaling 5 fingers.
The EgoGesture 3D dataset is a first-person perspective dataset with 24161 gesture samples recorded by 50 subjects across various environments, extracted from the original dataset EgoGesture through the MediaPipe [36] API. This dataset encompasses 83 types of static or dynamic gestures, primarily focusing on interactions with wearable devices. It complements the SHREC 2017 dataset, offering a broader environmental, accounting for lighting and background changes, as well as gestures performed with one hand and both hands, for validating gesture recognition methods.
Method Task0 Task 1 Task 2 Task 3 Task 4 Task 5 Task 6 Mean
G↑ G↑ IFM↓ G↑ IFM↓ G↑ IFM↓ G↑ IFM↓ G↑ IFM↓ G↑ IFM↓ G↑ IFM↓
Base [27] 78.1 60.4 13.5 18.9 63.3 9.3 82.4 8.0 84.3 5.9 88.5 5.9 88.3 18.1 70.0
Fine-tuning [27] 78.1 59.2 10.4 15.9 66.5 9.3 82.2 7.7 84.4 5.2 89.7 5.0 90.1 17.1 70.6
Feature extraction [27] 78.1 69.8 14.5 60.8 17.2 51.5 30.4 46.4 33.7 41.2 39.4 36.8 42.9 51.1 29.7
LwF [27] 78.1 69.1 0.1 43.0 27.3 18.8 67.3 11.3 78.7 6.5 87.5 6.3 87.7 25.8 58.1
LwF.MC [11] 78.1 36.8 9.8 24.2 41.3 17.0 64.6 12.8 73.3 10.2 78.3 9.7 80.2 18.4 57.9
DeepInversion [7] 78.1 68.1 14.1 44.3 32.2 24.7 59.2 16.2 71.2 11.6 78.8 10.0 81.0 29.1 56.1
ABD [11] 78.1 68.8 14.8 61.1 17.1 54.3 26.2 49.0 31.1 43.2 37.5 39.0 40.6 52.6 27.9
R-DFCIL [37] 78.1 70.3 3.0 61.4 4.1 53.2 14.9 46.1 17.8 39.2 35.0 35.2 36.3 50.9 18.5
BOAT-MI [8] 78.1 75.3 4.6 70.9 7.3 64.0 15.1 58.6 16.3 52.3 26.8 45.8 32.2 61.2 17.7
PGPFR (ours) 79.3 76.8 4.3 73.8 11.9 72.2 3.7 71.0 1.8 69.8 0.6 68.4 2.1 73.0 4.1
Oracle 75.8 75.8 75.8 75.8 75.8 75.8 75.8 75.8
TABLE I: Comparison with state-of-the-art approaches of class-incremental gesture recognition on the EgoGesture 3D dataset.
Method Task0 Task 1 Task 2 Task 3 Task 4 Task 5 Task 6 Mean
G↑ G↑ IFM↓ G↑ IFM↓ G↑ IFM↓ G↑ IFM↓ G↑ IFM↓ G↑ IFM↓ G↑ IFM↓
Base [27] 90.5 78.3 7.8 53.3 29.6 30.0 53.7 27.9 56.0 12.2 78.3 11.9 78.6 35.6 50.7
Fine tuning [27] 90.5 78.6 6.0 57.4 26.2 34.2 48.8 34.0 48.6 16.4 71.8 16.1 71.9 39.5 45.5
Feature extraction [27] 90.5 79.9 9.3 69.3 12.6 62.3 19.1 56.5 24.1 50.6 24.3 45.1 32.4 60.6 20.3
LwF [27] 90.5 79.8 9.3 64.1 20.2 31.9 50.2 29.1 53.1 13.1 76.8 11.5 79.3 38.2 48.2
LwF.MC [11] 90.5 56.0 10.9 32.6 39.5 21.8 58.5 19.3 61.6 16.7 53.3 16.1 45.4 27.1 44.9
DeepInversion [7] 90.5 79.9 4.3 65.9 14.9 53.1 29.5 49.5 32.8 34.2 47.7 32.1 49.7 52.5 29.8
ABD [11] 90.5 78.8 4.0 64.6 12.6 54.3 20.3 53.2 24.6 46.1 20.0 40.4 23.3 56.2 17.5
R-DFCIL [37] 90.5 78.7 3.3 65.5 4.5 54.4 20.5 49.8 26.9 41.5 25.0 38.6 33.3 54.8 18.9
BOAT-MI [8] 90.5 83.7 4.5 76.0 7.3 71.4 7.0 69.4 13.3 64.1 9.5 58.1 11.2 70.5 8.8
PGPFR (ours) 91.2 90.2 0.7 89.4 1.0 83.7 5.8 78.3 14.3 74.5 7.7 75.8 11.1 83.3 6.8
Oracle 89.4 89.4 89.4 89.4 89.4 89.4 89.4 89.4
TABLE II: Comparison with state-of-the-art approaches of class-incremental gesture recognition on the SHREC 2017 3D dataset.
IV-B Experimental Setup
The optimization strategy is Adaptive Moment Estimation (Adam) with a learning rate of 0.001. The batch size is 32. The initial training task covers 150 epochs, whereas each subsequent incremental training task comprises 100 epochs. An experiment is set up with a total of 7 incremental tasks. The temperature parameter $R$ of sharpening is set to 0.3, which can be adjusted with the variation of the dataset. The hyperparameter of the penalized term $\gamma$ is set to 1. The implementation is based on the PyTorch, and all experiments are performed on an NVIDIA RTX 4090 GPU with CUDA 12.0. The parameter count of our model is approximately 0.27M, which is light-weighted.
To assess the method, we utilize two types of test sets: one encompassing all visible categories and another consisting of current categories. For the SHREC 2017 3D dataset, after the model is trained on the initial incremental task (comprising 8 categories), we keep the backbone constant, while 1 new category is added to the visible classes during each subsequent incremental task. Similarly, for the EgoGesture 3D dataset, the backbone remains fixed once the model is trained on the first incremental class task (containing 59 categories), with 4 new categories being added during each subsequent incremental class task.
A critical issue in the development of incremental/continual learning systems lies in the disparity between the retention of information about previously learned tasks/classes and that of the newly introduced ones. As a result, a significant imbalance arises between the accuracy evaluated for current classes and the accuracy evaluated for all visible classes. We use the Instantaneous Forgetting Measure:
$IFM=\frac{|L-G|}{L+G}\times 100,$ (9)
proposed in [8] to measure the balance between our model’s retention of knowledge of past categories and its learning of knowledge of new categories. In particular, L represents the testing accuracy for new classes within the current task, while G (Global Accuracy) represents the testing accuracy for all visible classes up to the current task. The smaller the gap between L and G, the more balanced the model’s ability to retain knowledge of old classes and acquire knowledge of new classes. In the extreme case where the accuracy of either L or G is 0, the value of IFM reaches its maximum, indicating that the model’s capacity to balance these two abilities is severely limited. In the experimental tables, we primarily reported the evaluation indicators G (where a higher value is preferable) and IFM (where a lower value is desirable).
Method Task0 Task1 Task2 Task3 Task4 Task5 Task6 Mean
G↑ G↑ IFM↓ G↑ IFM↓ G↑ IFM↓ G↑ IFM↓ G↑ IFM↓ G↑ IFM↓ G↑ IFM↓
PGPFR (ours) 91.2 90.2 0.7 89.4 1.0 83.7 5.8 78.3 14.3 74.5 7.7 75.8 11.1 83.3 6.8
w/o PFGBP 91.2 89.0 2.9 86.8 4.0 80.6 3.9 76.5 7.7 73.7 5.9 75.0 12.5 81.8 6.2
w/o Sharpening 91.2 90.1 0.6 89.1 1.2 83.7 5.8 77.9 34.5 74.1 7.5 75.4 6.6 83.1 9.4
w/o Batch Proto 91.2 86.9 2.1 86.3 5.8 81.5 0.5 77.0 3.5 74.2 3.7 72.9 14.7 81.4 5.0
w/o $\mathcal{L}_{V}$ 91.2 90.6 1.0 85.5 3.2 75.8 3.5 70.4 7.7 66.8 6.2 60.1 20.6 77.2 7.0
w/o $\mathcal{L}_{V}$ w/ $\mathcal{L}_{Proto}$ 91.2 90.4 1.1 87.4 0.5 71.9 11.6 67.5 0.2 66.3 10.7 65.4 13.6 77.2 6.3
w/o $\mathcal{L}_{T}$ 91.2 90.1 0.3 89.2 0.4 82.8 10.8 77.6 11.1 74.8 10.7 75.5 8.5 83.0 7.0
Baseline 90.5 78.8 4.0 64.6 12.6 54.3 20.3 53.2 24.6 46.1 20.0 40.4 23.3 56.2 17.5
Baseline w/ CCRT 91.2 83.6 6.9 77.5 11.1 63.0 12.0 66.9 17.2 58.0 15.4 48.8 34.4 69.9 16.2
TABLE III: Ablation studies of the proposed method on the SHREC 2017 3D dataset. Our baseline is the ABD [11]. CCRT denotes continual classifier re-training. Sharpening and Batch Proto denote temperature sharpening of predicted logits of pseudo features and batch class prototypes in the PFGBP. For simplicity, PGPFR w/o $\mathcal{L}_{V}$ is shortened as w/o $\mathcal{L}_{V}$, etc.
IV-C Results of Gesture Recognition
Table I shows the comparison of our approach with state-of-the-art approaches on the EgoGesture 3D dataset. Tasks 0 to 6 represent the sequential stages of continuous data-free class incremental learning, while Mean denotes the average result across all seven tasks. It can be observed that our method achieves an average increase of 11.8% in Global Accuracy when compared to BOAT-MI [8], which is a recent state-of-the-art approach for data-free class-incremental gesture recognition. Furthermore, in terms of IFM, our method demonstrates an average decrease of 13.6. Particularly, for task 6, our approach outperforms BOAT-MI [8] by 31.1 in terms of IFM.
The results on the SHREC 2017 3D dataset are shown in Table II. Similarly, our approach significantly outperforms state-of-the-art methods. For the Global Accuracy, our approach beats BOAT-MI [8] by 12.8%. The results clearly demonstrate the remarkable adaptability of our approach in accommodating new classes while preserving the decision boundaries of previously learned ones.
(a) Accuracy of new classes.
(b) Accuracy of old classes
Figure 3: Comparison of accuracy of new and old classes among ablated approaches for each task on the SHREC 2017 3D dataset.
Figure 4: T-SNE visualization of the deep features extracted from the SHREC 2017 3D dataset. The experiemnts of class-incremental learning are conducted with 7 tasks. (a) depicts the features of the first 8 classes from PGPFR after training on task 0, (b) shows the features of the same first 8 classes from PGPFR but after training on task 6, (c) illustrates the features of all 14 classes from PGPFR after training on task 6. In all visualizations, data points are colored according to their corresponding ground-truth labels.
Figure 5: Visualization of gesture samples and corresponding classification results between the proposed method and BOAT-MI [8]. For a given sample, the gesture skeleton sequence is represented on the left, while the confidence scores predicted by different methods are displayed on the right.
IV-D Ablation Studies
The proposed PGPFR mainly comprises Pseudo Feature Generation with Batch Prototypes (PFGBP), Variational Prototype Replay for Old Classes (VPR), Truncated Cross-entropy for New Classes (TCE), and Classifier Re-Training (CCRT). The PFGBP module utilizes temperature sharpening in Eq.(4) and batch class prototypes in Eq.(1). The VPR module utilizes a penalty term in Eq.(6). Without this term, the regularization degenerates to PGT, as defined in Eq.(5). Detailed results of ablation studies are shown in Table III.
IV-D1 Effect of PFGBP, Batch Protytypes and Sharpening
As an important component of the proposed PGPFR, PFGBP plays a crucial role in the performance of the model. Without PFGBP, we notice a decline of 1.5% in terms of the mean Global Accuracy, while mean IFM shows a decline of 0.6. The batch class prototypes improves the global accuracy by 1.9% as it produces more diversified pseudo features. The temperature sharpening decreases the mean IFM by 2.6, indicating that it effectively prevents catastrophic forgetting by regulating the confidences of predicted probabilities associated with pseudo features.
IV-D2 Effect of VPR and Prototype Replay
Another important component is VPR, which is defined as $\mathcal{L}_{V}$. By eliminating the VPR, we observe a decline of 6.1% in terms of the mean Global Accuracy and an increase of 0.2 in terms of the mean IFM. It proves that VPR contributes more effectively to the model’s classification ability. By comparing the performance of w/o $\mathcal{L}_{V}$ to that of w/o $\mathcal{L}_{V}$ w/ $\mathcal{L}_{Proto}$, We observe that using only the original prototype classification loss results in minimal impact on the model’s performance, thereby further proving the effectiveness of the penalty term based on covariance matrices and weight vectors of the classifier.
IV-D3 Effect of TCE
The TCE loss, defined as $\mathcal{L}_{T}$, encourages the model to learn the semantic information of different classes rather than emphasizing the domain discrepancy between real data and pseudo features. We find that removing TCE results in inferior performance for both the mean Global Accuracy and IFM.
IV-D4 Effect of CCRT
we removed the CCRT strategy for baseline, meaning that the parameters of the feature extractor was no longer fixed. The results showed a significant drop in global accuracy after removing the CCRT strategy. This indicates that without CCRT, fine-tuning the model on new tasks causes the feature extractor to overly adapt to the new class samples, gradually overshadowing the representations of the old classes, leading to overfitting on the new classes. Therefore, the CCRT strategy is crucial for providing a general and stable feature representation, effectively preventing overfitting caused by small sample sizes or different data distributions.
IV-D5 Further Analysis of the Accuracies of New and Old Classes
By integrating these components, we achieve the highest Global Accuracy while maintaining a nearly minimal value of IFM. The results clearly demonstrate the effectiveness of these components in class-incremental gesture recognition.
We further analyze the roles of the components in Figure 3. Given that the accuracy for all visible classes and its disparity with the accuracy of the current classes in $\mathcal{D}^{t}$ are reported in Table III, we exhibit New Accuracy and Old Accuracy. New Accuracy denotes the accuracy of all visible classes excluding those from $\mathcal{D}^{0}$, while Old Accuracy denotes the accuracy of classes in the initial task $\mathcal{D}^{0}$. We find that, without the PFGBP and the VPR, the Old Accuracy decreases significantly during class-incremental learning, which further validates the roles of the proposed PFGBP and VPR in preventing catastrophic forgetting. Without the TCE, the Old Accuracy remains almost unchanged, as this module is solely applied to the new classes $\mathcal{D}^{t}$.
IV-D6 Visualizations
We utilize t-SNE [38] to reduce the dimensionality of the learned features, allowing us to visualize them and qualitatively assess the effectiveness of our proposed method. The deep features $h=\varphi(f(x))$ are projected using t-SNE and are color-coded according to their ground-truth labels in the resulting visualization. By comparing Figure 4(a) and Figure 4(b), we observe that the first 8 classes learned from Task 0 maintain clear and distinct boundaries in the test set for Task 0. Impressively, these boundaries remain well-defined even after six incremental tasks, highlighting the strong effectiveness of our method in mitigating catastrophic forgetting. Figure 4(c) displays the features of all 14 classes after learned on Task 6, using PGPFR. Recall SHREC 2017 3D is divided into 7 tasks, the visualization reveals a clear distinction between different classes.
We compare our method with the BOAT-MI [8] for hand gesture recognition of certain samples, as shown in Figure 5. Our method outperforms BOAT-MI in recognizing certain challenging samples. For instance, for the sample labeled Shake in Figure 5, our model accurately identifies it with a high confidence of 1.0, while BOAT-MI mistakenly classifies it as Swipe Left with a confidence of 0.7. Similarly, for the sample labeled Swipe Up, despite relatively lower confidence, our method correctly classifies it, whereas BOAT-MI wrongly predicts it as Expand. These results demonstrate the effectiveness of our method on these challenging gesture samples.
V Conclusion
In this work, we study class-incremental gesture recognition and introduce a Prototype-Guided Pseudo Feature Replay (PGPFR). This approach primarily consists of Pseudo Feature Generation with Batch Prototypes (PFGBP), Variational Prototype Replay (VPR), Truncated Cross-Entropy (TCE) and Continual Classifier Re-Training (CCRT). Extensive experiments on popular gesture recognition benchmarks demonstrate the effectiveness of the proposed PGPFR as well as the components. Specifically, PFGBP and VPR collaboratively tackle the issue of catastrophic forgetting of old knowledge when learning new classes. TCE mitigates the domain difference between real data and pseudo features and improves performance for new classes. We also discover that classifier re-training enhances the performance of class-incremental learning by potentially preventing overfitting to instances of new classes. We believe that it will considerably contribute to research in both data-free class-incremental learning and gesture recognition.